# Regular AdS3 black holes from regularized Gauss-Bonnet coupling

### v2

@@ 1 Introduction
Thanks to its relative simplicity, gravity in three dimensions (3d) serves as a useful theoretical laboratory. For example, 3d general relativity (GR) with a negative cosmological constant admits a black hole solution named as the Banados-Teitelboim-Zanelli (BTZ) black hole [1, 2], which is locally isomorphic to the AdS3 spacetime but still has all the important properties of a black hole. The geometry is characterized by the mass and the angular momentum, making it globally different than the AdS3 spacetime. It also has a well-defined event horizon and laws of black hole thermodynamics are satisfied. One of the most important challenges of a quantum theory of gravity is to provide a microscopic derivation of the black hole entropy. One motivation to study gravity in 3d is that it can be derived without any direct reference to string theory or supersymmetry [3] by employing AdS3/CFT2 correspondence (see [4] for a review).
@@
In this work, building upon our earlier work [14], we employ a bi-vector-tensor coupling obtained by a regularization of the Gauss-Bonnet (GB) invariant within the Weyl geometry and 3d find regular black hole solutions. The GB invariant is a topological term in 4d and identically vanishes in 3d. However, as shown in [15], a well-defined coupling that contributes to field equations in 4d can be obtained, and it gives rise to black hole solutions with a primary hair, i.e. an additional integration constant that modifies the geometry. By using the regularized GB coupling with two Weyl vectors, regular black hole solutions in 4d were obtained in [16]. In [14], we presented AdS3 black holes with primary hair by employing the 3d version of the regularized coupling. Now, in complete parallelism with the 4d story, we obtain regular AdS3 black hole solutions by considering two Weyl vectors.
@@ 3 Solution in [14] revisited
from which the following Noether charge follows
@@ 4 Regular black hole solutions
Note that this deformed coupling follows from the regularization $\lim_{d\to 3}\frac{\widetilde{\mathcal{G}}[B]-\kappa\widetilde{\mathcal{G}}[C]}{d-3}=\mathcal{G}[B]-\kappa\,\mathcal{G}[C]$.
By the method of [44] that we have outlined in the previous section, studying the solutions of the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory is a straightforward exercise. We now present regular black hole solutions and their charged versions by coupling the theories to Born–Infeld electrodynamics. In Appendix A, charged black hole solutions obtained by coupling to Maxwell’s electrodynamics are given. Due to the logarithmic form of the Coulomb potential in 3d, they possess a curvature singularity.
@@ 4.1 Undeformed theory ($\kappa=1$)
As mentioned before, after obtaining the equations of motion by extremizing reduced action in (4.4), analogous to (3.10), we choose
$b_{1}=\frac{b_{0}}{Nf},\quad\quad\quad\quad c_{1}=\frac{c_{0}}{Nf},$ (4.5)
Then, the consistency of the equations forces $N=1$ and one finds
@@
and the only other independent curvature invariants in 3d ($R^{\mu}_{\ \nu}R^{\nu}_{\ \mu}$ and $R^{\mu}_{\ \rho}R^{\rho}_{\ \nu}R^{\nu}_{\ \mu}$) [45, 46, 47] are also finite at the origin. Therefore, we conclude that it is a regular black hole solution with two primary hairs. Note that $q_{b}=q_{c}$ makes $B=C$ on-shell, and one obtains a stealth BTZ black hole solution.
@@ 4.2 Undeformed theory ($\kappa=1$) coupled to Born–Infeld electrodynamics
Using the vector components in (4.5, 4.6) together with the contribution from the Born–Infeld electrodynamics, the metric function is found as
@@
$R=\frac{3\left(2q_{b}^{2}+m\sqrt{\kappa}\,\ell^{2}\left|(\kappa-1)q_{b}\right|\right)}{2(\kappa-1)\ell^{2}q_{b}^{2}}+\mathcal{O}(r^{2}),$ (4.15)
and the other curvature invariants are well-defined at the origin, which show that it is a regular black hole as long as $q_{b}>q_{c}$. When $q_{b}=q_{c}$, we have the stealth black hole solution of Einstein–Born–Infeld theory [48].
@@ 4.3 Deformed theory ($\kappa\neq 1$)
The deformed theory can also be solved through the same procedure and the vector components take the same form in (4.5, 4.6). While a general solution for the metric function is possible to obtain, we prefer not to present it here due to its cumbersome form. As in 4d [16], by choosing a particular relation between the hair parameters, i.e. $q_{c}=q_{b}/\kappa$, we can find a simpler solution given by
@@
$f_{+}$ $=-\frac{r^{2}}{(\kappa-1)\ell^{2}}+r^{2}\Lambda_{0}+\mathcal{O}(\ell^{2}),$ (4.17)
$f_{-}$ $=-r^{2}\Lambda_{0}+\mathcal{O}(\ell^{2}).$ (4.18)
@@ 4.4 Deformed theory ($\kappa\neq 1$) coupled to Born–Infeld electrodynamics
The effective cosmological constant $\Lambda$ is the same as in (4.19). It is straightforward to show that this is also a regular black hole solution.
@@ Appendix A Coupling to Maxwell’s electrodynamics
The components of the Weyl vectors are given in (4.5, 4.6), and the metric functions are as follows
@@
From the $r\to 0$ limit, the behavior of the Ricci scalar is found as

### v3

# Regular AdS3 black holes from a regularized Gauss-Bonnet coupling

@@ 1 Introduction
Thanks to its relative simplicity, gravity in three dimensions (3d) serves as a useful theoretical laboratory. For example, 3d general relativity (GR) with a negative cosmological constant admits a black hole solution named as the Banados–Teitelboim–Zanelli (BTZ) black hole [1, 2], which is locally isomorphic to the AdS3 spacetime but still has all the important properties of a black hole. The geometry is characterized by the mass and the angular momentum, making it globally different than the AdS3 spacetime. It also has a well-defined event horizon and laws of black hole thermodynamics are satisfied. One of the most important challenges of a quantum theory of gravity is to provide a microscopic derivation of the black hole entropy. One motivation to study gravity in 3d is that it can be derived without any direct reference to string theory or supersymmetry [3] by employing AdS3/CFT2 correspondence (see [4] for a review).
@@
While the most natural candidates for admitting exact black hole solutions are higher-order pure gravity theories with second-order field equations for the metric, the most general theory with this property, the Lovelock gravity [14, 15, 16], is only defined in $d>4$. The reason is that the lowest order correction to the Einstein-Hilbert action, the Gauss-Bonnet (GB) invariant, is a topological term in 4d and identically vanishes in 3d. Recently, it was shown in [17] that a well-defined coupling that contributes to field equations in 4d can be obtained through a regularization of the GB invariant within the Weyl geometry.
The regularization is based on the fact that the GB invariant is non-dynamical in $d<5$ independently of the connection that is used. Therefore, evaluating the difference between the GB invariants for two different connections in a generic $d$ dimension, one obtains, up to a boundary term, a non-trivial Lagrangian density multiplied by (d3)(d4)(d-3)(d-4), which can be removed by scaling the coupling in $d$ dimensions. This way, two different regularized versions of the Gauss-Bonnet coupling can be obtained in 3d and 4d. Considering one Weyl connection and one Christoffel connection, we can obtain a vector-tensor theory of the generalized Proca class [18]. On the other hand, starting with two Weyl connections leads to a bi-vector-tensor coupling. Having second-order field equations, these vector-tensor theories possess black hole solutions with interesting properties.
In 4d, these two different possibilities were studied in [17, 19]. In the first case, with a single Weyl vector, one finds black hole solutions with a primary hair, i.e. an additional integration constant that modifies the geometry. In the second case, with two different Weyl vectors, regular black hole solutions with two primary hairs arise. The regularity of the resulting solutions is not guaranteed from the form of the Lagrangian density, as, for example, in the case of the regular point charge solution of Born-Infeld electrodynamics. However, the 4d case suggests a natural generalization of these solutions to 3d.
The regularized GB invariant in 3d with a single Weyl vector was constructed in [20] and the resulting black holes with a primary hair was given. Here, we investigate the possibility of having two Weyl vectors and construct a bi-vector-tensor coupling. In complete parallelism with the 4d story, we obtain regular AdS3 black hole solutions. As discussed in the last section, these black hole solutions are expected to exhibit interesting thermodynamical properties and the microscopic derivation of their entropy through the AdS3/CFT2 correspondence is particularly intriguing because of the existence of two primary hairs.
@@ 3 Solution in [20] revisited
where $\lambda$ is a constant parameter. The corresponding Noether charge is as follows
@@ 4 Regular black hole solutions
Note that this deformed coupling follows from the regularization
$\lim_{d\to 3}\frac{\widetilde{\mathcal{G}}[B]-\kappa\widetilde{\mathcal{G}}[C]+(\kappa-1)\mathcal{G}}{d-3}=\mathcal{G}[B]-\kappa\,\mathcal{G}[C].$ (4.5)
[…]
$\Xi[W]=$  $\,6r\omega_{0}^{4}-2fN^{2}\left[\omega_{0}^{2}(1+2r\omega_{1})-f^{2}N^{2}w_{1}^{2}(3+2r\omega_{1})\right]f^{\prime}$
$+2f^{2}N^{2}\left[f^{2}N^{2}\left(-r\omega_{1}^{4}+4\omega_{1}\omega_{1}^{\prime}+4r\omega_{1}^{2}\omega_{1}^{\prime}\right)-2\left(2r\omega_{1}\omega_{0}^{\prime}\omega_{0}+\omega_{0}^{2}(2\omega_{1}+r\omega_{1}^{2}+2r\omega_{1}^{\prime})\right)\right]$ (4.6)
$\Gamma[W]=$  $\,4r\omega_{0}^{4}-2f^{4}N^{3}\omega_{1}^{2}(3+2r\omega_{1})N^{\prime}-4f^{4}N^{4}\omega_{1}(1+r\omega_{1})(\omega_{1}^{2}-\omega_{1}^{\prime})$
$-4f^{2}N^{2}\omega_{0}(1+2r\omega_{1})\omega_{0}^{\prime}+2f^{2}N\omega_{0}^{2}\left((1+2r\omega_{1})N^{\prime}-2N(\omega_{1}+r\omega_{1}^{\prime})\right)$ (4.7)
$\Psi[W]=$ $-2r\omega_{0}^{2}+fN\left(N\left(1+2r\omega_{1}\right)f^{\prime}+2f\left(r\omega_{1}N^{\prime}+N\left(\omega_{1}+r\omega_{1}^{2}+r\omega_{1}^{\prime}\right)\right)\right)$ (4.8)
$\Phi[W]=$ $-rN\omega_{0}^{2}f^{\prime}+f^{2}N^{3}\omega_{1}(1+r\omega_{1})f^{\prime}+2f^{3}N^{2}\omega_{1}(1+2r\omega_{1})(N\omega_{1}+N^{\prime})$
$-2rf\omega_{0}\left(\omega_{0}N^{\prime}+N(\omega_{0}\omega_{1}-\omega_{0}^{\prime})\right)$ (4.9)
where
[…]
With these at hand, the equations of motion of the bi-vector theory can be written as
$\Psi[B]=0,\quad\quad\Psi[C]=0,\quad\quad\Phi[B]=0,\quad\quad\Phi[C]=0,$ (4.11)
$-f^{2}N^{4}(2r\Lambda_{0}+f^{\prime})+\ell^{2}(\Xi[B]-\kappa\Xi[C])=0,\quad\quad f^{3}N^{3}N^{\prime}+\ell^{2}(\Gamma[B]-\kappa\Gamma[C])=0$ (4.12)
We now present regular black hole solutions and their charged versions by coupling the theories to Born–Infeld electrodynamics. In Appendix A, charged black hole solutions obtained by coupling to Maxwell’s electrodynamics are given. Due to the logarithmic form of the Coulomb potential in 3d, they possess a curvature singularity.
@@ 4.1 Undeformed theory ($\kappa=1$)
Analogous to (3.10), we choose
$b_{1}=\frac{b_{0}}{Nf},\quad\quad\quad\quad c_{1}=\frac{c_{0}}{Nf}.$ (4.13)
Then, the consistency of the equations (4.11, 4.12) forces $N=1$, and one finds
@@
Figure 1: Figures showing the metric function and the corresponding curvature scalars for the undeformed theory. On the left, the plot of the metric function of the uncharged black hole in (4.15) is shown for a fixed set of parameters ($\kappa=1$, $m=1$, $\Lambda_{0}=-1$, and $\ell=0.5$) and for various values of the ratio $q_{b}/q_{c}$ (noted in the text boxes in the figure). The dashed line represents the BTZ black hole solution. On the right, metric function of the charged black hole in (4.21) is shown. The fixed parameters are the same with that of plot on the left, and the values of $q_{b}/q_{c}$ and $q_{c}$ varies as shown in the text boxes. The dotted line indicates the effect of a larger electric charge $q_{e}$ on the metric. In the insets, one can see the regularity of the curvature scalars explicitly.
@@
and the only other independent curvature invariants in 3d ($R^{\mu}_{\ \nu}R^{\nu}_{\ \mu}$ and $R^{\mu}_{\ \rho}R^{\rho}_{\ \nu}R^{\nu}_{\ \mu}$) [47, 48, 49] are also finite at the origin. For different sets of parameters, this solution is either a regular horizonless spacetime or a regular black hole with a double horizon. For a fixed parameter set, the metric function given in (4.15) has the behavior shown in Figure 1 (left), where we emphasize the effect of $q_{b}/q_{c}$ ratio on the horizon structure. While $q_{b}/q_{c}=1.05$ gives an extremal black hole solution, where the horizons coincide, any smaller and larger values $q_{b}/q_{c}$ lead double and horizonless cases respectively.
Note that $q_{b}=q_{c}$ makes $B_{\mu}=C_{\mu}$ on-shell. This is a case where the effective energy-momentum tensor vanishes on-shell and one obtains a stealth BTZ black hole solution, whose first example in the 3d literature was given in [50].
@@ 4.2 Undeformed theory ($\kappa=1$) coupled to Born–Infeld electrodynamics
Using the vector components in (4.13, 4.14) together with the contribution from the Born–Infeld electrodynamics, the metric function is found as
@@
$R=\frac{3[-q_{e}^{4}+2q_{e}^{2}(q_{b}+q_{c}+m)-8b^{2}\ell^{2}(q_{b}^{2}-q_{c}^{2})]}{4\ell^{2}q_{e}^{2}(q_{b}-q_{c})}+\mathcal{O}(r^{2})$ (4.23)
and the other curvature invariants are well-defined at the origin, which show that it is a regular black hole as long as $q_{b}>q_{c}$. In Figure 1 (right), the behavior of the metric function in (4.21) for different values of the electric charge $q_{e}$ can be seen. Considering the same fixed parameter set of parameters with that of the uncharged solution, presence of the electric charge shifts the critical $q_{b}/q_{c}$ ratio to larger values. For example, $q_{e}=1$ leads to an extremal black hole for $q_{b}/q_{c}=1.08$ instead of $1.05$. Increasing $q_{e}$ to 2, the metric function typically has larger values for all $r$.
When $q_{b}=q_{c}$, we have the black hole solution of Einstein–Born–Infeld theory [51] as a stealth solution.
@@ 4.3 Deformed theory ($\kappa\neq 1$)
The deformed theory can also be solved through the same procedure and the vector components take the same form in (4.13, 4.14). While a general solution for the metric function is possible to obtain, we prefer not to present it here due to its cumbersome form. As in 4d [19], by choosing a particular relation between the hair parameters, i.e. $q_{c}=q_{b}/\kappa$, we can find a simpler solution given by
@@
$f_{+}$ $=-\frac{r^{2}}{(\kappa-1)\ell^{2}}+\Lambda_{0}r^{2}+\mathcal{O}(\ell^{2}),$ (4.25)
$f_{-}$ $=-\Lambda_{0}r^{2}+\mathcal{O}(\ell^{2}).$ (4.26)
@@
Unlike the undeformed theory, this solution might also correspond to black holes with a single horizon in addition to horizonless geometries and black holes with double horizon. As an example, we consider a case with $m=1$, $\Lambda_{0}=-1$, $\ell=0.5$, and $q_{b}=1$ and study the effect of the deformation parameter $\kappa$ on the solution. We find the following behavior,
[…]
In Figure 2 (left), these different possibilities together with the regular behavior of the Ricci scalar are demonstrated.
Figure 2: Figures showing the metric function and the corresponding curvature scalars for the deformed theory. The same parameters that are used in Figure 1 are implemented here as well as $q_{b}=1$. On the left, the effect of the deformation parameter $\kappa$ on the metric function (4.24) can be seen, giving rise to different horizon structures. The dotted line is for the extremal case corresponding to $\kappa=8.52$. The dashed line represents the BTZ black hole solution. On the right, the metric function (4.30) of the charged black hole is shown for different $\kappa$ values, and $q_{e}=1$ and $q_{e}=2$. The presence of the electric charge $q_{e}$ increases the required $\kappa$ value for the extremal solution to 12.05.
@@ 4.4 Deformed theory ($\kappa\neq 1$) coupled to Born–Infeld electrodynamics
The effective cosmological constant $\Lambda$ is the same as in (4.27). It is straightforward to show that this is also a regular black hole solution. In Figure 2 (right) one can see that the overall behavior of the solution is quite similar to the uncharged case summarized in (4.29). A nonzero electric charge $q_{e}$ shifts the upper and lower values of the deformation parameter $\kappa$ generating different horizon structures towards larger values.
@@ Appendix A Coupling to Maxwell’s electrodynamics
The components of the Weyl vectors are given in (4.13, 4.14), and the metric functions are as follows
@@
Note that the relation between the hair parameters $q_{c}=\frac{q_{b}}{\kappa}$ was used in obtaining the latter solution as done in the main text. From the $r\to 0$ limit, the behavior of the Ricci scalar is found as